Two spherical objects have masses of 250 kg and 550 kg. Their centers are separated by a distance of 45 m. Find the gravitational attraction between them.




1
Expert's answer
2021-11-18T15:09:13-0500

"F = \frac{Gm_1m_2}{r^2} \\

F = \frac{6.67 \times 10^{-11} \times 250 \times 550}{45^2} = 4.53 \times 10^{-9} \;N"